OpenStudy (anonymous):

pleasee help verify that the intermediate value theorem applies to the indicated interval and find the value of c guaranteed by the theorem. f(x) = (x^2+x)/(x-1), [5/2,4], f(c) = 6

OpenStudy (anonymous):

\[f(x) = \frac{ x^2+x }{ x-1 }\]

OpenStudy (anonymous):

function is only discontinuous at 1, which is not in the interval theorem holds

OpenStudy (anonymous):

hmm... for the other questions, I plugged in c to x and set that number (in this case, 6) equal to the equation and solved it. but for this one, I can't

OpenStudy (anonymous):

\[f(x) = \frac{ x^2+x }{ x-1 }\] you need some numbers to get the second part

OpenStudy (anonymous):

oh here

OpenStudy (anonymous):

you need \(f(4)\) first

OpenStudy (anonymous):

[5/2 , 4] f(c) = 6

OpenStudy (anonymous):

the statement \(f(c)=6\) makes no sense

OpenStudy (anonymous):

well, it's given

OpenStudy (anonymous):

mvt says under the suitable conditions there is a \(c\in (a, b)\) where \[f'(c)=\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

oooh abort sorry wrong i am wrong

OpenStudy (anonymous):

oh I didn't learn that yet but okay.. what should I do?

OpenStudy (anonymous):

it thought it said mean value theorem, it says intermediate value theorem we have to start over sorry

OpenStudy (anonymous):

it's fine

OpenStudy (anonymous):

they way you show there must be a number in \([\frac{5}{2},4]\) with \[f(c)=6\] is to compute \(f(\frac{5}{2})\) and \(f(4)\)

OpenStudy (anonymous):

one number will be larger than 6, one number will be smaller than 6, so by the intermediate value theorem the function must be 6 somewhere in that interval

OpenStudy (anonymous):

since the function is continuous on that interval and cannot skip values

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry for the interruption but

OpenStudy (anonymous):

well, in class, I learned that c has to be replaced with x and have to set that equation to 6. when I get the x values, the values between 5/2 and 4 is the answer is this right?

OpenStudy (anonymous):

you job is to solve \[f(x)=6\] but that is for the second part of the question

OpenStudy (anonymous):

there are two parts 1) verify that the intermediate value theorem applies to the indicated interval

OpenStudy (anonymous):

the reason that you know \[f(x)=6\] for some \(x\) in the interval is because \(f(4)\) is larger than 6, and \(f(\frac{5}{2})\) is smaller than 6

OpenStudy (anonymous):

i think \(f(\frac{5}{2})=5\tfrac{5}{6}\) and \(f(4)=\frac{20}{3}\)

OpenStudy (anonymous):

how can 4 be larger than 6?

OpenStudy (anonymous):

k lets go slow

OpenStudy (anonymous):

it is not 4 that is larger than 6, it is \(f(4)=\frac{20}{3}\) that is larger than 6

OpenStudy (anonymous):

in other words \[f(\frac{5}{2})<6<f(4)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so since 6 lies between the two function values, there must be some \[c\in (\frac{5}{2},4)\] where \(f(c)=6\)

OpenStudy (anonymous):

now solving it is just a matter of solving \[\frac{x^2+x}{x-1}=6\]

OpenStudy (anonymous):

you would start with \[x^2+x=6(x-1)\] and solve the resulting quadratic equation

OpenStudy (anonymous):

ooh okay okay I finally get it :)

OpenStudy (anonymous):

Thank you so much for your kind explanation! :)

OpenStudy (anonymous):

yw